Reports 1-1 of 1 Clear search Modify search
OBS (General)
Tomaru - 1:09 Wednesday 15 April 2020 (14151) Print this report
Shift Report: 5pm-1am, 14th Apr. (volunteer)
Nakano-kun tried interferometer lock from this evening since ocean was going to be silent.
Around 11pm, interferometer was locked successfully.
However, interferometer is not still stable and lock-loss happened frequently.
Observation shift was took over to Ushiba-kun and Yuzurihara-kun.
Search Help